Q2-2
y (0) = 0, y (t) = 0 6) u - x?u+ dy=0 | u (0) = u (1) = 0 Q2: Find eigenvalues and eigenfunctions of the Sturm-Liouville s
Answer #1

2.2 u + du= 0, ula) = u(b) = 0 Case I :- dao u=o u=cat& c.fc₂> constant U(a) = 0 g carsh = 0 41b) –0 07b+shao ca-C,6=0 [c,sinub ula) = c, Casuar arsin ya o ulb)= c, casub + C₂ sin ub=0 x - x unta sinua C Casua sinub-c, Casub sinula = 0 → C [ sin (